8 years ago my father was 3 times as old as i shall be in next 5 years, when i was born my father was 32 years old how old am I?
Answer by ankor@dixie-net.com(22740)


let f = fathers age
let y = your age
:
Write an equation for each statement
:
8 years ago my father was 3 times as old as i shall be in next 5 years,
f - 8 = 3(y + 5)
f - 8 = 3y + 15
f = 3y + 15 + 8
f = 3y + 23
:
when i was born my father was 32 years
y = f - 32
Replace f with (3y+23) from the 1st statement
y = (3y+23) - 32
y = 3y - 9
9 = 3y - y
9 = 2y
y = 9/2
y = 4.5 yrs your age now